OpenStudy (anonymous):

Two end points of a line segment are (-12, -2) and (-6, -10). What are the coordinates of the point on the line through which its bisector passes? (-18, -12) (-9, -6) (-7, -8) (-11, -4)

OpenStudy (anonymous):

@cwrw238 can you help

OpenStudy (debbieg):

You need the midpoint of the line segment, right? do you know the formula for the midpoint of 2 points?

OpenStudy (debbieg):

http://www.purplemath.com/modules/midpoint.htm

OpenStudy (anonymous):

midpoint formula is \[\frac{ x1+x2 }{ 2 };\frac{ y1-y2 }{ 2 }\]

OpenStudy (debbieg):

Almost...\[\large \left( \frac{ x_{1}+x_{2} }{ 2 },\frac{ y_{1}+y_{2} }{ 2 } \right)\]

OpenStudy (anonymous):

alright i will try it

OpenStudy (anonymous):

@heather040200 @DebbieG (-9,-6) :) ?

OpenStudy (debbieg):

Yup :)

OpenStudy (anonymous):

yup I got that as well :)